Light from a LED enters through the pupil of the eye whose diameter is 6.1 mm. If the intensity is 2.3x10-11 W/m2, calculate the power of light that enters into the eye. If the LED is replaced with a He-Ne LASER ((=632.8 nm) calculate the power if 1.7x1015 photons/s are emitted. Assume that the laser beams diameter and pupil size are the same.


1
Expert's answer
2021-07-13T10:41:53-0400

a)

"P_1=IA=\frac{\pi Id^2}{4}=6.72\cdot 10^{-18}~W,"

b)

"P_2=\frac{vhc}{\lambda}=0.53~mW."